# Leveraging Large Language Models for Sarcastic Speech Annotation in Sarcasm Detection

### v1

@@ Abstract
Sarcasm fundamentally alters meaning through tone and context, yet detecting it in speech remains a challenge due to data scarcity. In addition, existing detection systems often rely on multimodal data, limiting their applicability in contexts where only speech is available. To address this, we propose an annotation pipeline that leverages large language models (LLMs) to generate a sarcasm dataset. Using a publicly available sarcasm-focused podcast, we employ GPT-4o and LLaMA 3 for initial sarcasm annotations, followed by human verification to resolve disagreements. We validate this approach by comparing annotation quality and detection performance on a publicly available sarcasm dataset using a collaborative gating architecture. Finally, we introduce PodSarc, a large-scale sarcastic speech dataset created through this pipeline. The detection model achieves a 73.63% F1 score, demonstrating the dataset’s potential as a benchmark for sarcasm detection research.
@@ 1 Introduction
Despite its importance, sarcasm detection in speech remains underexplored and is often constrained by small annotated datasets and challenges in generalization [1, 2, 3]. Publicly available datasets such as the Multimodal Sarcasm Detection Dataset (MUStARD) [4] and its extension, MUStARD++ [5] compiled from TV shows, include multimodal data and enable video-level multimodal sarcasm detection. However, these multimodal datasets are limited in size and scope, hindering the development of robust sarcasm detection models.
@@
In addition, most existing approaches to detecting sarcasm rely on multimodal feature fusion [2, 6, 7, 8]. However, in real-world applications, sarcasm often needs to be detected from speech alone, without the benefit of visual context. This challenges the development of sarcasm detection systems for audio-only scenarios like podcasts, radio broadcasts, and telephone conversations. Meanwhile, text-based sarcasm detection has progressed significantly in natural language processing (NLP). However, subtle cues in speech cannot be captured in this manner, since sarcasm is often conveyed not just through content, but also through auditory features such as intonation, pitch, pacing, and emphasis [9, 10]. More importantly, the scarcity of sarcastic utterances makes it difficult for detection models to learn fine-grained audio representations. As a result, most existing approaches rely on low-level acoustic features [5]. Thus, we urgently need large-scale annotated datasets specifically targeting sarcastic speech detection without visual cues.
@@
Recently, LLMs have shown promise in various NLP tasks, including emotion [11, 12], humor [13] and sarcasm understanding [14]. However, most advancements have been in text and image modalities [15]. This work leverages LLMs to facilitate sarcasm annotation of real-world “in-the-wild” speech data, specifically Overly Sarcastic Podcast (OSPod) [^1] https://overlysarcasticpodcast.transistor.fm, a series with a rich and diverse use of naturalistic and spontaneous sarcastic expressions. Human annotators are incorporated to resolve disagreements between LLM-based annotations.
@@
This work makes three key contributions: (1) We demonstrate the feasibility of using LLMs to facilitate sarcasm annotation, enabling the creation of a bi-modal sarcastic speech dataset with reduced reliance on human annotators. (2) We validate the effectiveness of our LLM-based annotation approach by comparing annotation performance and testing detection models on a publicly available dataset. (3) Using this pipeline, we create a large-scale high-quality sarcastic speech dataset that can be used for speech-based bi-modal sarcasm detection, with potential applications in real-world, speech-only environments.  [^2] Detailed prompts used for LLM annotations, as well as the annotated dataset before and after human verification, will be made available for research purposes upon request.
@@ 2 Related Work
Datasets for sarcasm detection The detection of sarcasm in speech has been limited by the availability of annotated datasets. General emotion detection datasets, such as MELD [16] and IEMOCAP [17], while rich in their inclusion of emotional speech, do not specifically capture the subtleties of sarcasm. While sarcasm can overlap with emotions like anger, joy, or surprise, it possesses distinct features that emotion-focused annotations fail to capture. Existing sarcasm detection datasets like MUStARD [4] and MUStARD++ [5], although instrumental in advancing multimodal sarcasm detection, primarily focus on incorporating visual and textual cues alongside speech. Since some sarcastic labels in these datasets rely on visual context (including facial expressions, gestures), audio-only models may show reduced performance. Therefore, their applicability to real-world speech-only environments is limited.
Automatic speech data collection Gathering large-scale, high-quality training data with appropriate labels has always been a critical aspect of speech technology development. Over decades, the speech community has invested substantial effort in collecting and annotating speech data with segmentation, transcription, and speaker labels, significantly advancing speech recognition and synthesis technologies. To automate the data collection process, automatic preprocessing frameworks that generate segmentation, speaker labels, and transcriptions for in-the-wild speech while removing noise, reverberation, and overlapping speech have been proposed [18, 19].
LLMs for data annotation Model training heavily depends on datasets with high-quality annotations. Given their success in annotation tasks, LLMs are increasingly being explored as data annotators for scenarios with limited or well-defined label sets [20, 21, 22]. Recent studies have explored the affective capabilities of LLMs, demonstrating their ability to infer emotions in given contexts and respond with emotional support in dialogues [15]. Specifically, LLMs like GPT have shown promise in identifying emotions and intent from text [11, 23]. LLMs have also been evaluated in zero- or few-shot emotion recognition tasks and have shown good performance [11, 24, 25]. Compared with general emotion recognition tasks, sarcasm detection using LLMs remains underexplored, likely due to its greater reliance on nuanced understanding beyond general emotional cues.
@@ 3.1 Automatic data collection and processing
As illustrated in Figure 1, the raw audio from OSPod was systematically processed to extract high-quality speech segments at utterance level and generate corresponding transcripts using Emilia-Pipe [19], an open-source preprocessing pipeline designed to convert in-the-wild speech data into structured datasets with precise annotations (e.g., timestamps, speaker information). In total, 30 episodes were processed, yielding 11,024 segments. Each segment was aligned with its transcript to create a robust bimodal dataset for annotation. To enhance data quality, the processed dataset was carefully checked to remove noise, irrelevant content, and unclear segments.
@@ 3.2 Sarcasm annotation using LLMs
For the processed dataset, we used two LLMs as annotators, GPT-4o [^3] https://chat.openai.com and LLaMA 3 [26], to perform sarcasm label annotations. These models were selected for their state-of-the-art performance in understanding contextual and nuanced language. For each annotator, we provide a carefully designed prompt tailored to guide the LLM in accurately identifying sarcasm (Figure 1). The prompts include structured raw texts, highlighted cues to focus on, examples of sarcastic and non-sarcastic expressions, and an annotation output template. The annotation process was carried out with both GPT-4o and LLaMA 3 independently labeling each segment. This dual-annotator approach enabled us to compare the outputs and identify samples of agreement and disagreement between the two models.
To validate the effectiveness of LLM-based annotation, we compared the LLMs’ annotations with the gold labels of the MUStARD++ dataset. Table 1 presents the annotation performance of GPT-4o and two versions of LLaMA 3 on the MUStARD++ dataset. Among the models, GPT-4o achieved the highest performance, with a Macro-averaged F1 (Macro-F1) of 67.47% and a Unweighted Average Recall (UAR) of 67.12%, indicating its strong capability in sarcasm annotation. The LLaMA 3 models performed comparably, with LLaMA 3-70B slightly outperforming the 8B variant (63.59% vs. 61.52% in F1 score). [^4] We will use LLaMA 3-70B only for the following experiments. These results suggest that all evaluated LLMs exhibit reasonable proficiency in sarcasm detection.
@@ 3.3 Human verification
With the LLMs’ annotation capabilities confirmed, we proceeded to annotate the OSPod dataset. While both models showed strong performance, they produced conflicting labels for 2,884 of the 11,024 utterances. To resolve these discrepancies and ensure high-quality annotations, two PhD students specializing in phonetics manually reviewed these cases.
Since cases where LLMs produced conflicting labels were highly ambiguous, inter-annotator agreement was calculated with a Kappa score of 0.58. The differences were reconciled by a third annotator. Figure 1 illustrates how the discrepancies were addressed during the human verification phase. Annotators first reviewed the explanations provided by the LLMs. Since prosodic cues were missing during the LLM annotation phase, the human review placed particular emphasis on prosody alongside textual features. For example, in the utterance: “Usually we’re separated by oceans and stars, but now we’re in the same space, so, uh.” The phrase “so, uh.” features a playful tone and hesitation. Combined with the dramatic phrasing of “oceans and stars” and the conversational context of a first-time in-person recording, this utterance was labeled as sarcastic.
@@ 4.1 Experimental setup
We follow MUStARD++ for feature extraction and sarcasm detection [5]. MUStARD++ introduces a collaborative gating architecture that enables multimodal feature fusion, followed by classification through fully connected layers. While the original implementation integrates text, audio, and visual modalities [^8] https://github.com/cfiltnlp/MUStARD_Plus_Plus, our work focuses on the text and audio modalities. Below, we briefly describe the extraction process for each modality.
Text modality (T) We encode the text using BART [27] Large model with $d_{t}=1024$ and use the mean of the last four transformer layer representations to get a unique embedding representation for each utterance.
@@
We conduct two detection experiments using MUStARD++ and PodSarc. First, we evaluate our annotation pipeline on MUStARD++ to validate that LLM-generated labels can produce functional detection models compared to human annotations. This validation establishes the viability of our approach before applying it to create PodSarc. Second, we evaluate detection performance on PodSarc itself to demonstrate the dataset’s utility. MUStARD++ is split into training and testing sets at an 8:2 ratio (962:240). For PodSarc, we use the entire dataset for training, reserving three manually annotated episodes for testing with a sarcasm-to-non-sarcasm ratio of 243:504. For all training runs, we perform hyperparameter tuning with dropout values in [0.2, 0.3, 0.4], learning rates in [0.001, 0.0001], batch sizes in [32, 64, 128], shared embedding sizes in [1024, 2048], and projection embedding sizes in [256, 1024].
@@ 4.3 Performance on PodSarc
Compared to the detection results on MUStARD++, models trained on LLM-annotated labels on PodSarc show better performance across all settings (e.g., A$+$T setting: 71.47% vs. 61.8%). This difference may be due to dataset composition. MUStARD++ incorporates three modalities (text, audio, and visual), with some sarcastic utterances manually annotated based on visual cues—an aspect absent in our LLM-based annotation method and PodSarc. Since our detection models cannot leverage visual context, their performance on MUStARD++ is lower. In contrast, PodSarc consists solely of text and audio-based sarcasm, aligning closely with the features available to the models.

### v2

@@ Abstract
Sarcasm fundamentally alters meaning through tone and context, yet detecting it in speech remains a challenge due to data scarcity. In addition, existing detection systems often rely on multimodal data, limiting their applicability in contexts where only speech is available. To address this, we propose an annotation pipeline that leverages large language models (LLMs) to generate a sarcasm dataset. Using a publicly available sarcasm-focused podcast, we employ GPT-4o and LLaMA 3 for initial sarcasm annotations, followed by human verification to resolve disagreements. We validate this approach by comparing annotation quality and detection performance on a publicly available sarcasm dataset using a collaborative gating architecture. Finally, we introduce PodSarc, a large-scale sarcastic speech dataset created through this pipeline. The detection model achieves a 73.63% F1 score, demonstrating the dataset's potential as a benchmark for sarcasm detection research.
@@ 1 Introduction
Despite its importance, sarcasm detection in speech remains underexplored and is often constrained by small annotated datasets and challenges in generalization [rakov2013sure, gao22f_interspeech, li2023sarcasticspeech]. Publicly available datasets such as the Multimodal Sarcasm Detection Dataset (MUStARD) [castro-etal-2019-towards] and its extension, MUStARD++ [ray-etal-2022-multimodal] compiled from TV shows, include multimodal data and enable video-level multimodal sarcasm detection. However, these multimodal datasets are limited in size and scope, hindering the development of robust sarcasm detection models.
@@
In addition, most existing approaches to detecting sarcasm rely on multimodal feature fusion [gao22f_interspeech, cai-etal-2019-multi, gao2024amused, raghuvanshi2025intra]. However, in real-world applications, sarcasm often needs to be detected from speech alone, without the benefit of visual context. This challenges the development of sarcasm detection systems for audio-only scenarios like podcasts, radio broadcasts, and telephone conversations. Meanwhile, text-based sarcasm detection has progressed significantly in natural language processing (NLP). However, subtle cues in speech cannot be captured in this manner, since sarcasm is often conveyed not just through content, but also through auditory features such as intonation, pitch, pacing, and emphasis [loevenbruck2013prosodic, li2024functional]. More importantly, the scarcity of sarcastic utterances makes it difficult for detection models to learn fine-grained audio representations. As a result, most existing approaches rely on low-level acoustic features [ray-etal-2022-multimodal]. Thus, we urgently need large-scale annotated datasets specifically targeting sarcastic speech detection without visual cues.
@@
Recently, LLMs have shown promise in various NLP tasks, including emotion [santoso2024large, zhang2024refashioning], humor [chen2023can] and sarcasm understanding [zhang2024sarcasmbench]. However, most advancements have been in text and image modalities [zhao2023chatgpt]. This work leverages LLMs to facilitate sarcasm annotation of real-world ``in-the-wild" speech data, specifically Overly Sarcastic Podcast (OSPod) [^1] https://overlysarcasticpodcast.transistor.fm, a series with a rich and diverse use of naturalistic and spontaneous sarcastic expressions. Human annotators are incorporated to resolve disagreements between LLM-based annotations.
@@
This work makes three key contributions: (1) We demonstrate the feasibility of using LLMs to facilitate sarcasm annotation, enabling the creation of a bi-modal sarcastic speech dataset with reduced reliance on human annotators. (2) We validate the effectiveness of our LLM-based annotation approach by comparing annotation performance and testing detection models on a publicly available dataset. (3) Using this pipeline, we create a large-scale high-quality sarcastic speech dataset that can be used for speech-based bi-modal sarcasm detection, with potential applications in real-world, speech-only environments.  [^2] Detailed prompts used for LLM annotations, as well as the annotated dataset before and after human verification are available at https://github.com/Abel1802/PodSarc.
@@ 2 Related Work
Datasets for sarcasm detection The detection of sarcasm in speech has been limited by the availability of annotated datasets. General emotion detection datasets, such as MELD [poria2019meld] and IEMOCAP [busso2008iemocap], while rich in their inclusion of emotional speech, do not specifically capture the subtleties of sarcasm. While sarcasm can overlap with emotions like anger, joy, or surprise, it possesses distinct features that emotion-focused annotations fail to capture. Existing sarcasm detection datasets like MUStARD [castro-etal-2019-towards] and MUStARD++ [ray-etal-2022-multimodal], although instrumental in advancing multimodal sarcasm detection, primarily focus on incorporating visual and textual cues alongside speech. Since some sarcastic labels in these datasets rely on visual context (including facial expressions, gestures), audio-only models may show reduced performance. Therefore, their applicability to real-world speech-only environments is limited.
Automatic speech data collection Gathering large-scale, high-quality training data with appropriate labels has always been a critical aspect of speech technology development. Over decades, the speech community has invested substantial effort in collecting and annotating speech data with segmentation, transcription, and speaker labels, significantly advancing speech recognition and synthesis technologies. To automate the data collection process, automatic preprocessing frameworks that generate segmentation, speaker labels, and transcriptions for in-the-wild speech while removing noise, reverberation, and overlapping speech have been proposed [yu2024autoprep, he2024emilia].
LLMs for data annotation Model training heavily depends on datasets with high-quality annotations. Given their success in annotation tasks, LLMs are increasingly being explored as data annotators for scenarios with limited or well-defined label sets [ding2023gpt, moller2024parrot, tan2024large]. Recent studies have explored the affective capabilities of LLMs, demonstrating their ability to infer emotions in given contexts and respond with emotional support in dialogues [zhao2023chatgpt]. Specifically, LLMs like GPT have shown promise in identifying emotions and intent from text [santoso2024large, niu2024text]. LLMs have also been evaluated in zero- or few-shot emotion recognition tasks and have shown good performance [santoso2024large, feng2023affect, sun2023text]. Compared with general emotion recognition tasks, sarcasm detection using LLMs remains underexplored, likely due to its greater reliance on nuanced understanding beyond general emotional cues.
@@ 3.1 Automatic data collection and processing
As illustrated in Figure 1, the raw audio from OSPod was systematically processed to extract high-quality speech segments at utterance level and generate corresponding transcripts using Emilia-Pipe [he2024emilia], an open-source preprocessing pipeline designed to convert in-the-wild speech data into structured datasets with precise annotations (e.g., timestamps, speaker information). In total, 30 episodes were processed, yielding 11,024 segments. Each segment was aligned with its transcript to create a robust bimodal dataset for annotation. To enhance data quality, the processed dataset was carefully checked to remove noise, irrelevant content, and unclear segments.
@@ 3.2 Sarcasm annotation using LLMs
For the processed dataset, we used two LLMs as annotators, GPT-4o [^3] https://chat.openai.com and LLaMA 3 [touvron2023LLaMA], to perform sarcasm label annotations. These models were selected for their state-of-the-art performance in understanding contextual and nuanced language. For each annotator, we provide a carefully designed prompt tailored to guide the LLM in accurately identifying sarcasm (Figure 1). The prompts include structured raw texts, highlighted cues to focus on, examples of sarcastic and non-sarcastic expressions, and an annotation output template. The annotation process was carried out with both GPT-4o and LLaMA 3 independently labeling each segment. This dual-annotator approach enabled us to compare the outputs and identify samples of agreement and disagreement between the two models.
To validate the effectiveness of LLM-based annotation, we compared the LLMs' annotations with the gold labels of the MUStARD++ dataset. Table 1 presents the annotation performance of GPT-4o and two versions of LLaMA 3 on the MUStARD++ dataset. Among the models, GPT-4o achieved the highest performance, with a Macro-averaged F1 (Macro-F1) of 67.47% and a Unweighted Average Recall (UAR) of 67.12%, indicating its strong capability in sarcasm annotation. The LLaMA 3 models performed comparably, with LLaMA 3-70B slightly outperforming the 8B variant (63.59% vs. 61.52% in F1 score). [^4] We will use LLaMA 3-70B only for the following experiments. These results suggest that all evaluated LLMs exhibit reasonable proficiency in sarcasm detection.
@@ 3.3 Human verification
With the LLMs' annotation capabilities confirmed, we proceeded to annotate the OSPod dataset. While both models showed strong performance, they produced conflicting labels for 2,884 of the 11,024 utterances. To resolve these discrepancies and ensure high-quality annotations, two PhD students specializing in phonetics manually reviewed these cases.
Since cases where LLMs produced conflicting labels were highly ambiguous, inter-annotator agreement was calculated with a Kappa score of 0.58. The differences were reconciled by a third annotator. Figure 1 illustrates how the discrepancies were addressed during the human verification phase. Annotators first reviewed the explanations provided by the LLMs. Since prosodic cues were missing during the LLM annotation phase, the human review placed particular emphasis on prosody alongside textual features. For example, in the utterance: ``Usually we're separated by oceans and stars, but now we're in the same space, so, uh." The phrase ``so, uh." features a playful tone and hesitation. Combined with the dramatic phrasing of ``oceans and stars" and the conversational context of a first-time in-person recording, this utterance was labeled as sarcastic.
@@ 4.1 Experimental setup
We follow MUStARD++ for feature extraction and sarcasm detection [ray-etal-2022-multimodal]. MUStARD++ introduces a collaborative gating architecture that enables multimodal feature fusion, followed by classification through fully connected layers. While the original implementation integrates text, audio, and visual modalities [^8] https://github.com/cfiltnlp/MUStARD_Plus_Plus, our work focuses on the text and audio modalities. Below, we briefly describe the extraction process for each modality.
Text modality (T) We encode the text using BART [lewis2019bart] Large model with $d_{t}=1024$ and use the mean of the last four transformer layer representations to get a unique embedding representation for each utterance.
@@
We conduct two detection experiments using MUStARD++ and PodSarc. First, we evaluate our annotation pipeline on MUStARD++ to validate that LLM-generated labels can produce functional detection models compared to human annotations. This validation establishes the viability of our approach before applying it to create PodSarc. Second, we evaluate detection performance on PodSarc itself to demonstrate the dataset's utility. MUStARD++ is split into training and testing sets at an 8:2 ratio (962:240). For PodSarc, we use the entire dataset for training, reserving three manually annotated episodes for testing with a sarcasm-to-non-sarcasm ratio of 243:504. For all training runs, we perform hyperparameter tuning with dropout values in [0.2, 0.3, 0.4], learning rates in [0.001, 0.0001], batch sizes in [32, 64, 128], shared embedding sizes in [1024, 2048], and projection embedding sizes in [256, 1024].
@@ 4.3 Performance on PodSarc
Compared to the detection results on MUStARD++, models trained on LLM-annotated labels on PodSarc show better performance across all settings (e.g., A++T setting: 71.47% vs. 61.8%). This difference may be due to dataset composition. MUStARD++ incorporates three modalities (text, audio, and visual), with some sarcastic utterances manually annotated based on visual cues—an aspect absent in our LLM-based annotation method and PodSarc. Since our detection models cannot leverage visual context, their performance on MUStARD++ is lower. In contrast, PodSarc consists solely of text and audio-based sarcasm, aligning closely with the features available to the models.